# Towards Trustworthy Dermatology MLLMs: A Benchmark and Multimodal Evaluator for Diagnostic Narratives

### v1

@@ 1 Introduction
Dermatologic diseases are prevalent across populations and impose a persistent clinical burden (Karimkhani et al., 2017; Urban et al., 2021). The widespread availability of smartphone and clinical imaging has created a rich visual substrate for computational analysis (Bourkas et al., 2023; Farr et al., 2021; Ouellette & Rao, 2022). Recent advances in multimodal LLMs (Zhou et al., 2024; Nasir, 2025) have made it possible to transform a single dermatology image into a full diagnostic narrative (Yan et al., 2025; Zeng et al., 2025; Lin et al., 2025) that describes visible findings, articulates reasoning, and proposes differential diagnoses (Sellergren et al., 2025; Liu et al., 2023; Zhu et al., 2023; Hurst et al., 2024; Li et al., 2023a; Singhal et al., 2025). This capability is attractive for scalable decision support (Yang et al., 2025; Han et al., 2023; Güneş et al., 2025; Gabashvili, 2023; Pillai et al., 2024; Liu et al., 2024), yet it is also fragile (Asgari et al., 2025; Nakaura et al., 2025; Zack et al., 2024). A narrative that is fluent but clinically unsound can mislead users and create safety risks (Asgari et al., 2025; Chustecki, 2024). Responsible deployment therefore requires rigorous and clinically grounded evaluation of generated texts (Aljamaan et al., 2024).
@@
We then establish a systematic evaluation framework that comprises DermBench and DermEval, as shown in Figure 1. DermBench pairs fixed images with physician approved exemplars and uses an LLM-based judge to compare candidate narratives against certified references. It then assigns scores along clinically grounded dimensions, enabling controlled evaluation of the diagnostic generation capabilities of multimodal models. DermEval is a multimodal evaluator that maps an image and a narrative to scalar scores and a structured critique. At inference time, the model requires no reference text. Given only an image and a diagnostic narrative, it produces an evaluation narrative and per case scores that enable routine assessment. Training proceeds in two stages. In the first stage, the model learns a canonical, machine‑parsable format via token‑level cross‑entropy. In the second stage, it aligns predicted scores with physician ratings using a reinforcement objective with an exponential moving‑average baseline, restricting policy gradients to generated tokens. We adopt six dimensions that capture essential requirements, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision.
Our key contributions are:
@@ 2.1 General and Domain-Specific Multimodal Models
In the medical domain, LLaVA-Med adapts biomedical figure–caption corpora (Li et al., 2023b). HuatuoGPT-Vision scales medical visual alignment with improvements on health benchmarks (Chen et al., 2024). Med-PaLM M explores generalist multimodal modeling across clinical modalities (Tu et al., 2024). Domain-specific dermatology foundation models further push scale: PanDerm achieves state-of-the-art performance on 28 benchmarks including clinician reader studies (Yan et al., 2025), while DermINO leverages 432K hybrid-pretrained skin images for tasks such as classification, captioning, and fairness evaluation (Xu et al., 2025). Yet across both general-purpose and domain-specific models, the absence of dermatology CoT supervision limits their ability to produce reliable, structured reasoning for clinical workflows.
@@ 3 Methodology
In this section we present an integrated pipeline that spans data construction, benchmarking, and evaluator learning. We first describe the dataset making process in Section 3.1, which builds paired image and reasoning data through a dual stream procedure that uses structured prompts for caption generation and hierarchical clinical reasoning, followed by expert curation that produces certified references and evaluation rationales for each image and diagnostic narrative. We then define DermBench in Section 3.2, which pairs each image with its certified reference and employs an LLM judge to score candidate narratives on six dimensions, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision, thereby establishing the official metrics for reporting. Finally we introduce DermEval in Section 3.3, a LLaVA based evaluator that maps an image and a narrative to six scalar scores and a structured evaluation. Training proceeds in two stages. Stage one learns a canonical and parsable evaluation format with token level cross entropy. Stage two aligns emitted scores to physician scores using score oriented REINFORCE with an exponential moving average baseline while sending policy gradients only to generated tokens. Section 3.4 specifies the unified objective that combines the text loss and the reinforcement loss with fixed weights and summarizes the inference time prompting and score extraction protocol.
@@ 3.1 Dataset Construction
We construct the dataset through two coordinated streams, as shown in Figure 2. In the high-quality stream we first select a large set of dermatology images with diverse and balanced categories. For each image we use Gemini 2.5 Pro (Comanici et al., 2025)to prompt a vision language model to produce a caption that reports only observable findings such as anatomic location, lesion morphology, and surface characteristics while forbidding diagnostic speculation. The captioning prompt is: “Please briefly describe the physical location of the affected skin and the observable medical features of the skin condition. Do not make any differential diagnosis. Start your response with ‘This image shows …’.” A second LLM then receives the caption together with the ground truth label and is instructed to simulate clinical reasoning. The hierarchical reasoning prompt is: “Simulate expert hierarchical reasoning for dermatologic diagnosis using the provided caption and the ground truth {DISEASE_NAME}. Begin with high level categorization, progressively refine to specific diseases and pathological features, and conclude with a coherent diagnostic judgment.” The model reasons stepwise and presents the disease name only in the final sentence. We normalize the output into a hierarchical chain of thought that states coarse disease families, then intermediate descriptors, and finally the specific diagnosis. Board certified dermatologists rate every text on six axes from 0 to 5, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. If any dimension receives a score below 5, clinicians revise the text until all dimensions are rated 5, after which the recorded scores are fixed at 5.
@@
High score texts constitute the benchmark introduced in Section 3.2. For each image, its diagnostic narrative, and the associated six‑dimension scores, we additionally elicit an evaluation rationale using the following instruction: “Given the dermatology image, the generated diagnostic narrative, and the six numeric scores for Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision, produce a concise justification for each dimension. Structure the output as six titled sections that match the dimension names. In each section restate the score, cite concrete evidence from the narrative or observable findings that supports the score, and end with one actionable suggestion for improvement. Do not propose a new diagnosis and do not alter the scores.” The resulting rationale is saved as the evaluation text label. All image–diagnostic text–evaluation text triples, including both high and low score cases, are used to train DermEval in Section 3.3.
@@ 3.2 DermBench Construction and LLM Scoring
We construct DermBench from the images and certified high-quality reference texts defined in the previous subsection, as shown in Figure 3. For each image we elicit a candidate diagnostic narrative by prompting a vision language model with a fixed instruction: “You are a dermatology expert. Analyze the image by describing key features such as morphology, color, distribution, and surface changes. Interpret the findings with clinical reasoning. Provide two to three differential diagnoses. Justify the top diagnosis. Conclude with ‘The answer is {DISEASE_NAME}’ where {DISEASE_NAME} is the disease name.” The resulting narrative is treated as the candidate text. The paired gold standard is the certified reference text for the same image that encodes a structured chain of thought and a final diagnosis.
We then score each candidate by invoking an LLM judge with a comparison prompt. The judge receives the candidate text and the gold reference text and is instructed to assign a score from 0 to 5 on six dimensions, specifically Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. The comparison instruction states: “Given the two passages above, where the first is our generated diagnostic text and the second is the gold standard reference, compare them and assign our generated text a score from 0 to 5 for Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. Use 0 for the lowest and 5 for the highest.” The six scores constitute the official DermBench metrics.
@@ 3.3 DermEval Training with Score Oriented REINFORCE with EMA Baseline
We finetune a LLaVA model into DermEval that maps an image and a diagnostic text to six scalar scores and a structured evaluation text, as shown in Fig 4. The six dimensions are Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. Physician scores on these six dimensions are normalized to the interval from zero to five. Training proceeds in two stages. Stage 1 teaches the model to produce a canonical and parsable evaluation format. Stage 2 aligns the scores emitted in the generated evaluation text with physician scores through reinforcement learning.
In Stage 1 the model receives an image $I$ and a diagnostic text $d$. The target output is a templated evaluation text $y^{*}$ that contains six explicit score fields. The model is optimized with token level cross entropy to obtain stable structured outputs that can be reliably parsed in the next stage
@@
where $x$ denotes the prompted input that is composed by the conversation template and the image sentinel with $(I,d)$. The variable $y$ is the predicted evaluation text and $y^{*}$ is the target template.
In Stage 2 we introduce Score Oriented REINFORCE with exponential moving average (EMA) Baseline. For each instance we construct the same prompt as in inference and generate a single evaluation text $\hat{y}$ without gradients. An external LLM parses $\hat{y}$ to extract the six scores $\hat{\mathbf{s}}\in[0,5]^{6}$, with the prompt for the external LLM parser shown in Figure 4. Let $\mathbf{s}^{*}\in[0,5]^{6}$ be the physician scores, and let $\mathcal{I}_{y}$ be the set of indices that are successfully parsed with size $K_{y}$. The numeric alignment reward is the negative mean squared error on valid dimensions
@@ 4.1 Alignment Tests with Expert Annotations
To verify the alignment of DermBench and DermEval with expert judgment, we selected nine representative models spanning three categories, namely general purpose multimodal models including Gemini-2.5-Flash, GPT-4o mini, and GPT-o4 mini, reasoning-focused multimodal models including Vision-R1, InternVL3, LLaVA-o1, and Qwen2.5-VL, and medical domain models including HuatuoGPT-Vision and LLaVA-Med-7B. We evaluated 500 images sampled from DermNet that were not used for DermEval training to avoid overfitting. For each image and for each model we generated a diagnostic narrative under the standardized instruction protocol defined in Section 3, yielding 4500 narratives in total. Dermatologists then assigned scores on six clinically grounded dimensions for every narrative. In addition, for each image we produced a perfect reference narrative following the procedure in Section 3.1. We subsequently evaluated all model generated narratives with DermBench using the certified reference and with DermEval without any reference. Finally we computed the per dimension errors of DermBench and DermEval relative to physician scores. All scores are rounded to the nearest integer on the 0 to 5 scale in order to align with physician ratings. Results are reported in Table 1.
@@
Evaluator Acc Safe MedG Cover Reason Desc
DermBench 0.251 0.314 0.369 0.456 0.412 0.377
DermEval 0.117 0.230 0.176 0.152 0.236 0.147
Both DermBench and DermEval track physician judgments closely, with all errors below 0.5 on the 0 to 5 scale. The macro average MAE is 0.36 for DermBench and 0.18 for DermEval, which indicates low discrepancy overall and a consistent advantage for DermEval. DermEval achieves lower MAE on every dimension, with the largest reductions on Clinical Coverage and Description Precision, while DermBench remains competitive as a reference-anchored comparator. These findings support DermBench for controlled comparisons against certified references and DermEval for case level scoring without references, and show that the two evaluators provide reliable and complementary mechanisms for measuring diagnostic quality. Additionally, we present examples in the Supplementary Material showing that the scores from DermBench and DermEval closely match physician ratings.

### v2

@@ 1 Introduction
Dermatologic diseases are prevalent across populations and impose a persistent clinical burden (Karimkhani et al., 2017; Urban et al., 2021). The widespread availability of smartphone and clinical imaging has created a rich visual substrate for computational analysis (Bourkas et al., 2023; Farr et al., 2021; Ouellette & Rao, 2022). Recent advances in multimodal LLMs (Zhou et al., 2024; Nasir, 2025) have made it possible to transform a single dermatology image into a full diagnostic narrative (Yan et al., 2025; Zeng et al., 2025; Lin et al., 2025) that describes visible findings, articulates reasoning, and proposes differential diagnoses (Sellergren et al., 2025; Liu et al., 2023; Zhu et al., 2023; Hurst et al., 2024; Li et al., 2023; Singhal et al., 2025). This capability is attractive for scalable decision support (Yang et al., 2025; Han et al., 2023; Güneş et al., 2025; Gabashvili, 2023; Pillai et al., 2024; Liu et al., 2024b), yet it is also fragile (Asgari et al., 2025; Nakaura et al., 2025; Zack et al., 2024). A narrative that is fluent but clinically unsound can mislead users and create safety risks (Asgari et al., 2025; Chustecki, 2024). Responsible deployment therefore requires rigorous and clinically grounded evaluation of generated texts (Aljamaan et al., 2024).
@@
We then establish a systematic evaluation framework that comprises DermBench and DermEval, as shown in Fig. 1. DermBench pairs fixed images with physician approved exemplars and uses an LLM-based judge to compare candidate narratives against certified references. It then assigns scores along clinically grounded dimensions, enabling controlled evaluation of the diagnostic generation capabilities of multimodal models. DermEval is a multimodal evaluator that maps an image and a narrative to scalar scores and a structured critique. At inference time, the model requires no reference text. Given only an image and a diagnostic narrative, it produces an evaluation narrative and per case scores that enable routine assessment. Training proceeds in two stages. In the first stage, the model learns a canonical, machine parsable format via token level cross entropy. In the second stage, it aligns predicted scores with physician ratings using a reinforcement objective with an exponential moving average baseline, restricting policy gradients to generated tokens. We adopt six dimensions that capture essential requirements, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. More detailed definitions of these six metrics are provided in Sec. A.1. Our key contributions are:
@@ 2.1 General and Domain-Specific Multimodal Models
In the medical domain, LLaVA-Med adapts biomedical figure–caption corpora (Li et al., 2023). HuatuoGPT-Vision scales medical visual alignment with improvements on health benchmarks (Chen et al., 2024). Med-PaLM M explores generalist multimodal modeling across clinical modalities (Tu et al., 2024). Domain-specific dermatology foundation models further push scale: PanDerm achieves state-of-the-art performance on 28 benchmarks including clinician reader studies (Yan et al., 2025), while DermINO leverages 432K hybrid-pretrained skin images for tasks such as classification, captioning, and fairness evaluation (Xu et al., 2025). Yet across both general-purpose and domain-specific models, the absence of dermatology CoT supervision limits their ability to produce reliable, structured reasoning for clinical workflows.
@@ 2.2 Evaluators and Model-as-Judge Approaches
2.3 Chain-of-thought Reasoning in Medical Contexts
Chain-of-thought reasoning trains models to generate explicit intermediate steps rather than only final answers and improves performance on complex language tasks (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022; Zhou et al., 2022; Zhang et al., 2022). In clinical applications, chain-of-thought supervision helps large models follow plausible diagnostic reasoning, achieve higher accuracy on medical question answering and improve perceived interpretability (Nachane et al., 2024; Jeon & Kim, 2025). MedCoT introduces hierarchical expert guided medical chains of thought aligned with domain knowledge (Liu et al., 2024a), while ReasonMed and MedReason provide large scale medical reasoning trajectories for training and analysis (Sun et al., 2025; Wu et al., 2025). These developments motivate our focus on dermatology specific diagnostic narratives and on evaluators that score chains of thought along clinically meaningful dimensions.
@@ 3 Methodology
In this section we present an integrated pipeline that spans data construction, benchmarking, and evaluator learning. Sec. 3.1 describes the dataset making process, which builds paired image and reasoning data through a dual stream procedure. Sec. 3.2 then defines DermBench, which pairs each image with its certified reference and employs an LLM judge to score candidate narratives on six dimensions, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision, thereby establishing the official metrics for reporting. DermBench is the first dermatology benchmark that evaluates diagnostic narratives instead of only categorical labels, uses a six dimension rubric that explicitly targets clinical trustworthiness, and includes certified reference narratives curated by dermatologists for every image. Finally we introduce DermEval in Sec. 3.3, a LLaVA based evaluator that maps an image and a narrative to six scalar scores and a structured evaluation. DermEval is the first multimodal evaluator trained directly on physician scores and it enables reference free evaluation that conditions jointly on the image and the diagnostic narrative. Sec. 3.4 specifies the unified objective that combines the text loss and the reinforcement loss with fixed weights and summarizes the inference time prompting and score extraction protocol.
@@ 3.1 Dataset Construction
We construct the dataset through two coordinated streams, as shown in Fig. 2. In the high-quality stream we first select a large set of dermatology images with diverse and balanced categories. For each image we use Gemini 2.5 Pro (Comanici et al., 2025) to prompt a vision language model to produce a caption that reports only observable findings such as anatomic location, lesion morphology, and surface characteristics while forbidding diagnostic speculation. The captioning prompt is: “Please briefly describe the physical location of the affected skin and the observable medical features of the skin condition. Do not make any differential diagnosis. Start your response with ‘This image shows …’.” A second LLM then receives the caption together with the ground truth label and is instructed to simulate clinical reasoning. The hierarchical reasoning prompt is: “Simulate expert hierarchical reasoning for dermatologic diagnosis using the provided caption and the ground truth {DISEASE_NAME}. Begin with high level categorization, progressively refine to specific diseases and pathological features, and conclude with a coherent diagnostic judgment.” The model reasons stepwise and presents the disease name only in the final sentence. We normalize the output into a hierarchical chain of thought that states coarse disease families, then intermediate descriptors, and finally the specific diagnosis. Board certified dermatologists rate every text on six axes from 0 to 5, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. If any dimension receives a score below 5, clinicians revise the text until all dimensions are rated 5, after which the recorded scores are fixed at 5. All images processed by closed-source models such as Gemini are sourced from DermNet and comply with institutional policies on public, de-identified medical datasets.
@@
High score texts constitute the benchmark introduced in Sec. 3.2. For each image, its diagnostic narrative, and the associated six-dimension scores, we additionally elicit an evaluation rationale using the following instruction: “Given the dermatology image, the generated diagnostic narrative, and the six numeric scores for Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision, produce a concise justification for each dimension. Structure the output as six titled sections that match the dimension names. In each section restate the score, cite concrete evidence from the narrative or observable findings that supports the score, and end with one actionable suggestion for improvement. Do not propose a new diagnosis and do not alter the scores.” The resulting rationale is saved as the evaluation text label. All image–diagnostic text–evaluation text triples, including both high and low score cases, are used to train DermEval in Sec. 3.3.
Two dermatologists served as raters and co-authors. One has more than four years of clinical, teaching, and research experience with multiple dermatology publications. The other is the Chief Dermatologist and Department Director in the same institution with more than twenty years of clinical and academic experience and membership in national dermatology committees. They first scored all narratives across six dimensions on their own and curated the high-quality reference narratives. Although inter-rater reliability was not formally quantified due to the small rater pool, all scores were mutually visible and both raters jointly reviewed and reconciled them, discussing any disagreement on a case until a unified decision was reached.
@@
Table 1: Comparison of DermBench and DermEval with representative dermatology benchmarks. Task abbreviations: Cls = classification, Seg = segmentation, Cap = captioning, Risk = risk prediction, DNE = diagnostic narrative evaluation.
Benchmark Task Modality
Free-text
support
Clinician
reference
Number of
metrics
Automated
evaluator
ISIC challenges Cls, Seg Image $\times$ $\times$ 1 $\times$
Derm7pt Cls Image+Tabular $\times$ $\times$ 1 $\times$
PAD-UFES-20 Cls Image+Tabular $\times$ $\times$ 1 $\times$
PanDerm Cls, Seg, Risk Image $\times$ $\times$ 3 $\times$
DermINO Cls, Cap, Seg Image $\checkmark$ $\times$ 3 $\times$
DermBench (ours) DNE Image+Text $\checkmark$ $\checkmark$ 6 $\checkmark$
DermEval (ours) DNE Image+Text $\checkmark$ $\checkmark$ 6 $\checkmark$
@@ 3.2 DermBench Construction and LLM Scoring
We construct DermBench from the images and certified high-quality reference texts defined in the previous subsection, as shown in Fig. 3. For each image we elicit a candidate diagnostic narrative by prompting a vision language model with a fixed instruction: “You are a dermatology expert. Analyze the image by describing key features such as morphology, color, distribution, and surface changes. Interpret the findings with clinical reasoning. Provide two to three differential diagnoses. Justify the top diagnosis. Conclude with ‘The answer is {DISEASE_NAME}’ where {DISEASE_NAME} is the disease name.” The resulting narrative is treated as the candidate text. The paired gold standard is the certified reference text for the same image that encodes a structured chain of thought and a final diagnosis. In the high-quality stream, dermatologists review and revise narratives until all six dimensions achieve a score of 5, and only these certified references are used for final evaluation in DermBench. As summarized in Tab. 1, unlike prior dermatology benchmarks that mainly focus on image-level tasks with a few aggregate metrics, DermBench and DermEval target diagnostic narrative evaluation on image plus text, use clinician-certified reference narratives, and score outputs along six explicit clinical dimensions.
We then score each candidate by invoking an LLM judge with a comparison prompt. The judge model in the DermBench evaluation pipeline is DeepSeek-R1. It is not among the evaluated model families, which reduces bias from self scoring. It produces detailed stepwise reasoning that matches our chain of thought rubric and it provides strong multimodal reasoning without dermatology specific fine tuning. The judge receives the candidate text and the gold reference text and is instructed to assign a score from 0 to 5 on six dimensions, specifically Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. The comparison instruction states: “Given the two passages above, where the first is our generated diagnostic text and the second is the gold standard reference, compare them and assign our generated text a score from 0 to 5 for Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. Use 0 for the lowest and 5 for the highest.” The six scores constitute the official DermBench metrics.
@@ 3.3 DermEval Training with Score Oriented REINFORCE with EMA Baseline
We finetune a LLaVA model into DermEval that maps an image and a diagnostic text to six scalar scores and a structured evaluation text, as shown in Fig. 4. The six dimensions are Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. Physician scores on these six dimensions are normalized to the interval from zero to five. Training proceeds in two stages. Stage 1 teaches the model to produce a canonical and parsable evaluation format. Stage 2 aligns the scores emitted in the generated evaluation text with physician scores through reinforcement learning.
In Stage 1 the model receives an image $I$ and a diagnostic text $d$. The target output is a templated evaluation text $y^{*}$ that contains six explicit score fields. The model is optimized with token level cross entropy to obtain stable structured outputs that can be reliably parsed in the next stage.
@@
In Stage 2 we introduce Score Oriented REINFORCE with exponential moving average (EMA) Baseline. For each instance we construct the same prompt as in inference and generate a single evaluation text $\hat{y}$ without gradients. An external LLM parses $\hat{y}$ to extract the six scores $\hat{\mathbf{s}}\in[0,5]^{6}$, with the prompt for the external LLM parser shown in Fig. 4. Let $\mathbf{s}^{*}\in[0,5]^{6}$ be the physician scores, and let $\mathcal{I}_{y}$ be the set of indices that are successfully parsed with size $K_{y}$. The numeric alignment reward is the negative mean squared error on valid dimensions
@@ 4.1 Alignment Tests with Expert Annotations
To verify the alignment of DermBench and DermEval with expert judgment, we selected nine representative models spanning three categories, namely general purpose multimodal models including Gemini 2.5 Flash, GPT-4o-mini, and GPT-o4-mini, reasoning-focused multimodal models including Vision-R1, InternVL3, LLaVA-o1, and Qwen2.5-VL, and medical domain models including HuatuoGPT-Vision and LLaVA-Med-7B. We evaluated 500 images sampled from DermNet that are completely disjoint from all training splits, including DermEval training and reference construction, and are used only for this alignment study. For each image and for each model we generated a diagnostic narrative under the standardized instruction protocol defined in Sec. 3, yielding 4500 narratives in total. Dermatologists then assigned scores on six clinically grounded dimensions for every narrative. In addition, for each image we produced a perfect reference narrative following the procedure in Sec. 3.1. We subsequently evaluated all model generated narratives with DermBench using the certified reference and with DermEval without any reference. Finally we computed the per dimension errors of DermBench and DermEval relative to physician scores. All scores are rounded to the nearest integer on the 0 to 5 scale in order to align with physician ratings. Results are reported in Tab. 2.
DermBench and DermEval yield highly consistent scores across all six metrics, with typical differences below 0.05 for the same model and metric. This tight agreement means that the reference free DermEval preserves the preference structure induced by the reference anchored DermBench rather than reshaping it. In both evaluators the same models occupy the top and bottom positions across all six metrics, which indicates that DermEval has internalized physician like scoring behavior and can serve as a reliable judge for case level assessment without requiring gold reference narratives.
@@
Evaluator Acc Safe MedG Cover Reason Desc Avg
DermBench 0.251 0.314 0.369 0.456 0.412 0.377 0.363
DermEval 0.117 0.230 0.176 0.152 0.236 0.147 0.178